OpenStudy (anonymous):

3.The shorter leg of a 30°- 60°- 90° right triangle is 12.5 inches. How long are the longer leg and the hypotenuse?

OpenStudy (anonymous):

that would be 25 degrees

OpenStudy (anonymous):

12.5/sin30

OpenStudy (anonymous):

I mean the hypotenuse is 25

OpenStudy (anonymous):

hypotenuse is 12.5*2=25 longer = sqrt(25^2-12.5^2)=21.6 (pytago) or because it is the half uniform triangle longer 25*sqrt(3)/2=21.6

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

and the other side would be 21.65 cos30*25.

OpenStudy (anonymous):

what is

jhonyy9 (jhonyy9):

because to the smaler angle provided is the smaler side --yes ? so than the shorter leg is 12,5 in than this is provide to angle with 30 degree - so do you know how many is sin30=? and sin60=? or cos20=? or cos60=?

jhonyy9 (jhonyy9):

do you know this trigonometric table ?

OpenStudy (anonymous):

no im doing pythagoream thoerum

jhonyy9 (jhonyy9):

yes but in this case you dont know just one side what sign that here you not can using the pythagorean teorem here

OpenStudy (anonymous):

ok

jhonyy9 (jhonyy9):

so how do you think if you know that to the shorter side is oposite angle is 30 degree and you know that sin of an angle is equale the opposite side divided by hypothenuse so from this can you calcule anything ,different side ?

jhonyy9 (jhonyy9):

yes i havemissed that you know again that sin30=1/2

jhonyy9 (jhonyy9):

robetoby do you understand it ?

jhonyy9 (jhonyy9):

pardon for your name

jhonyy9 (jhonyy9):

come on robey i think you will can solve it sure

jhonyy9 (jhonyy9):

so than sin 30 = shorther leg divide hypothenuse so 1/2 = 12,5/hypothenuse from this the hypothenuse =12,5*2=25 ao and now you can us pythagorean theorem so 25^2 =12,5^2 +longer side ^2 from this the longer side will be equal sqrt(25^2 -12,5^2) =sqrt(625-156,25)=sqrt(468,75) = =21,65

OpenStudy (anonymous):

ok